LHCb Daily Report, Wednesday 16 August 2017, 10:30

Shift Leader Data Manager GEOC
Yesterday morning Kristof De Bruyn Alessandro Bertolin Raja Nandakumar
Yesterday afternoon Frederic Teubert Meriem Boubdir
Yesterday night Paul Seyfert David Gerick
Today morning Matthew Charles Alessandro Bertolin Raja Nandakumar
Today afternoon Frederic Teubert Christopher Burr
Today night Paul Seyfert Emanuele Michielin

Luminosity and inefficiencies

Fill Delivered Logged Inefficiency HV off Velo not IN DAQ not running Dead time
6089 5694.311 nb-1 5364.843 nb-1 5.79 % 0.03 % 0.39 % 1.21 % 4.23 %
6086 2390.391 nb-1 2271.671 nb-1 4.97 % 0.09 % 0.25 % 0.40 % 4.26 %
Total 0.000 nb-1 0.000 nb-1 0.00 % 0.00 % 0.00 % 0.00 % 0.00 %

Latest events

Shift Summary entries in the logbook

15 August 14:30 Kristof De Bruyn LHC:
- 2.45 fill with 1356 bunches, dumped again by losses in 16L2
- Now: short access

LHCb:
- smooth running during stable beams, checks concerning /daqarea in background
(Although the app was enabled, LHC did take a very long time to reach our target lumi/mu)
- Access:
- Heinrich went down to check up on the DSS sensor warning for one of the supply bottle
- Beat is testing the storage system
- Karol has VELO in local to test new software

To be done:
- Make sure everybody returns their subsystems in time
15 August 22:20 Frederic Teubert LHC:
fill 6087 (with 1500 bunches) lost at ramp, just before flat top reached due to losses in 16L2.
Since then problems with BLMs at point 6 have prevented any further attempt to fill the machine (with less bunches, 1300).

LHCb:
Beat has been investigating the problem with daqarea.
Heinrich/Laurent have understood (not yet solved) the DSS alarm on gas supply bottle. The alarm has been cleared.
Karol has been working on the VELO LV.

To be done:
Take good data when LHC sorts the BLM issue.
16 August 06:34 Paul Seyfert LHC:
- physics with 1356 bunches for most of the shift

LHCb:
- problem with RICH1 LV at the begin of the fill, couldn't get run control running w/o excluding RICH1. Cured
itself while piquet was working on it https://lblogbook.cern.ch/Shift/117335
- manual run change because of OT errors https://lblogbook.cern.ch/Shift/117345
- Muon Alignment error message on the alarm screen at the moment https://lblogbook.cern.ch/Shift/117352 to be
fixed during day time
- TMUDAQ errors appear to cure themselves https://lblogbook.cern.ch/Shift/117353

To be done:
- take data

Run Chief : Paula Collins and Yuri Guz

VELO Piquet : Chris Parkes

Name Report
16 Aug 08:03:23 Chris Parkes
Pedestal drift observed on sensor 76 over the last three weeks is also related to lower LV voltages and
currents. A message has been introduced to appear to the shift leader message box if the LV reduces.

The error message appeared for the first time at the end of fill this morning.

ST Piquet : Eliseo Perez Trigo

Name Report
16 August 09:15 Eliseo Perez Trigo Nothing to report.

OT Piquet :

Name Report
16 August 09:03 Bartlomiej Rachwal All OK

RICH Piquet : Iwan Thomas Smith

Name Report
16 August 08:24 Iwan Smith Minor issue with one L0 board refusing to power on last night. After a few minutes of debugging, it managed to
turn on, and no physics time was lost.

Calo Piquet : Olivier Deschamps

Name Report
16 August 10:25 Olivier Deschamps All fine

Muon Piquet : Nikolai Bondar

Name Report
16 August 09:57 Oleg Maev Nothing to report.

HERSCHEL Piquet : Paolo Gandini

L0 Piquet : Olivier Deschamps

Name Report
16 August 10:25 Olivier Deschamps all fine

HLT Piquet : Alex Pearce

HLT1 disk occupancy HLT2 output rate

Alignment Piquet : Silvia Borghi

Name Report
16 August 09:52 Silvia Borghi Rich mirror alignment updated due to change of magnet polarity.
Rich alignment automatic update switched on again.

Online Piquet : Paolo Durante

Name Report
16 August 10:07 Paolo Durante /daqarea was temporarily brought down to check for data integrity errors, none were found.
Ran a stress test in the afternoon during EOF_CALIB and the filesystem did not crash, still investigating.

Data Quality Piquet : Michael Wilkinson

Production :

Name Report
15 August 18:05 Raja Nandakumar Trello/JIRA ticket:
https://lblogbook.cern.ch/Operations/27589
https://ggus.eu/?mode=ticket_info&ticket;_id=130018
https://ggus.eu/?mode=ticket_info&ticket;_id=130072 - solved now